each of the 100 students at a music school play at least one instrument, string, woodwind or brass. 70 play a string instrument, 49 woodwinds and 49 brasses. 20 plays both string and woodwind, 25 play string and brass, and 35 plays both woodwind and brass. How many play all three types of instrument?
Answer by ikleyn(52776)

.
We have the set A of 100 students and three its subsets S (string; 70 elements), W (wooding; 49 elements), and B (brasses; 49 elements).

Let us denote SW the intersection of S and W for brevity; denote SB the intersection of S and B; and denote
WB the intersection of W and B.

Also let us denote SWB the intersection of all the three subsets S, W and B.


We know that the subset S consists of 70 elements, the subset W consists of 49 elements and the subset B consists of 49 elements.

We are also given that SW consists of 20 elements, SB consists of 25 elements, and WB consists of 35 elements.

The question is: how many elements are in SWB?


For any given subset X, let us use the symbol |X| for the number of elements in X.


Then there is a remarkable equality, which connects the number of elements in the set A; in its subsets S, W and B;
in their intersections SW, SB and WB; and in the subset SWB:

|A| = |S| + |W| + |B| - |SW| - |SB| - |WB| + |SWB|.     (1)

under the condition that the union of subsets S, W and B covers the entire A:  A = S U W U B.


Now look how this equality will help us to solve our problem.
Simply substitute the known values into this equality. You will get

100 = 70 + 49 + 49 - 20 - 25 - 35 + |SWB|.                 (2)

|SWB| is the only unknown in this equation, and you can easily find it by isolating.

|SWB| = 100 - (70+49+49) + (20+25+35) = 12.


Answer.  12 students play all 3 instruments.

------------
See the lessons
    - Counting elements in sub-sets of a given finite set
    - Advanced problems on counting elements in sub-sets of a given finite set
in this site.